OpenStudy (anonymous):

James bought juice for his class party. The label says that there are 925 millimeters of juice in each container. Write a formula that helps james determine the servings of juice in each container. y= number of servings of juice z=number of millimeters per serving

OpenStudy (anonymous):

If 1 container = 925 mm juice And z = mm / serving then 1/z = serving / mm So we can obtain the number of servings in 925 mm by taking 925 * 1/z = 925 mm * 1 serving / mm = 925/z servings Since this is what the variable 'y' is supposed to represent, let y = 925/z